# M4-SAR: A Multi-Resolution, Multi-Polarization, Multi-Scene, Multi-Source Dataset and Benchmark for Optical-SAR Fusion Object Detection

### v1

@@ 4.2 Quantitative Evaluation
R-FCOS [49] ICCV’19 $\checkmark$ 31.9 17.6ms 32.8 79.8 29.7 67.0 78.9 54.8 57.2 26.4 30.2
R-ATSS [74] CVPR’20 $\checkmark$ 36.0 19.1ms 33.2 80.0 29.3 66.6 78.3 54.8 57.0 27.1 30.0
YOLOv5 [21] arXiv’20 $\checkmark$ 8.10 7.20ms 61.4 88.0 56.6 76.4 91.3 91.8 77.6 59.8 52.9
Roi-Trans [39] arXiv’21 $\checkmark$ 273 49.3ms 17.2 80.7 9.1 55.2 80.5 25.6 44.2 27.6 27.1
O-RepPoint [25] CVPR’22 $\checkmark$ 36.6 16.1ms 63.3 90.2 55.8 81.7 88.2 90.1 77.1 58.3 53.9
RTMDet [38] arXiv’22 $\checkmark$ 8.86 11.1ms 34.0 79.8 30.3 61.1 79.1 55.9 56.7 26.5 29.2
YOLOv8 [21] arXiv’23 $\checkmark$ 10.0 6.70ms 63.6 88.6 57.9 80.0 91.3 92.5 79.0 61.6 54.4
PSC [66] CVPR’23 $\checkmark$ 31.9 18.3ms 30.7 79.7 25.9 57.4 78.1 49.4 53.5 26.2 28.5
KFIoU [65] ICLR’23 $\checkmark$ - 66.2ms 16.9 79.4 9.1 50.9 80.4 25.2 43.6 19.4 22.4
YOLOv6 [22] ICLR’24 $\checkmark$ 16.0 6.10ms 57.6 82.1 52.7 67.6 89.4 90.4 73.3 64.9 48.8
YOLOv9 [56] ECCV’24 $\checkmark$ 6.40 14.4ms 61.9 87.7 55.3 74.4 90.0 92.4 76.9 58.9 52.5
YOLOv10 [53] NeurIPS’24 $\checkmark$ 7.50 7.80ms 65.2 89.9 58.8 82.3 90.8 93.5 80.1 62.8 55.7
YOLOv11 [21] arXiv’24 $\checkmark$ 9.74 6.50ms 64.3 90.0 57.8 80.7 90.9 91.5 79.2 60.5 54.3
LSKNet [27] IJCV’24 $\checkmark$ 21.8 20.4ms 63.3 90.2 55.8 81.7 88.2 90.1 77.1 58.3 53.9
R-FCOS [49] ICCV’19 $\checkmark$ 31.9 17.6ms 23.9 80.1 27.9 60.1 75.8 43.5 51.9 26.4 28.3
R-ATSS [74] CVPR’20 $\checkmark$ 36.0 19.1ms 25.3 80.3 26.8 62.3 77.0 43.9 52.6 24.9 27.8
YOLOv5 [21] arXiv’20 $\checkmark$ 8.10 7.20ms 40.5 85.3 41.0 64.3 85.5 89.9 67.7 44.0 41.3
Roi-Trans [39] arXiv’21 $\checkmark$ 273 49.3ms 9.1 80.8 9.1 35.6 79.0 18.0 38.6 25.2 24.3
O-RepPoint [25] CVPR’22 $\checkmark$ 36.6 16.1ms 26.2 78.6 21.9 53.7 59.9 48.8 48.2 13.9 20.4
RTMDet [38] arXiv’22 $\checkmark$ 8.86 11.1ms 16.2 71.0 21.1 39.3 68.9 39.6 42.7 20.9 22.3
YOLOv8 [21] arXiv’23 $\checkmark$ 10.0 6.70ms 44.9 85.8 42.5 67.9 87.5 89.2 69.6 46.1 43.2
PSC [66] CVPR’23 $\checkmark$ 31.9 18.3ms 22.3 80.0 20.6 51.9 75.7 39.4 48.3 24.5 26.5
KFIoU [65] ICLR’23 $\checkmark$ - 66.2ms 9.1 72.1 9.1 35.3 78.0 22.7 37.7 17.4 19.8
YOLOv6 [22] ICLR’24 $\checkmark$ 16.0 6.10ms 36.5 79.4 32.8 53.5 85.7 86.7 62.4 37.6 36.5
YOLOv9 [56] ECCV’24 $\checkmark$ 6.40 14.4ms 43.9 84.8 38.4 61.5 87.9 88.8 67.6 44.2 41.4
YOLOv10 [53] NeurIPS’24 $\checkmark$ 7.50 7.80ms 49.8 87.0 45.5 73.4 87.9 90.4 72.3 50.8 46.2
YOLOv11 [21] arXiv’24 $\checkmark$ 9.74 6.50ms 46.9 87.2 44.6 71.1 88.8 89.2 71.3 48.6 44.7
LSKNet [27] IJCV’24 $\checkmark$ 21.8 20.4ms 33.4 80.2 28.4 69.3 76.8 51.2 56.6 27.1 30.4
CFT [42] PR’22 $\checkmark$ $\checkmark$ 53.8 40.6ms 75.8 92.5 61.3 91.6 90.3 96.3 84.6 68.9 59.9
CLANet [18] ACMMM’23 $\checkmark$ $\checkmark$ 48.2 29.1ms 74.8 92.2 60.7 91.3 91.6 97.2 84.6 68.5 59.6
CSSA [2] CVPR’23 $\checkmark$ $\checkmark$ 13.5 12.3ms 73.3 91.7 59.3 88.9 91.6 95.8 83.4 66.4 58.0
CMADet [45] TIV’24 $\checkmark$ $\checkmark$ 41.5 46.7ms 70.9 90.7 52.0 86.4 91.7 97.1 81.5 63.5 55.7
ICAFusion [44] PR’24 $\checkmark$ $\checkmark$ 29.0 23.6ms 74.7 91.9 60.9 91.0 91.8 96.7 84.5 67.3 58.8
MMIDet [69] TCSVT’24 $\checkmark$ $\checkmark$ 53.8 41.9ms 74.9 92.6 61.1 91.7 91.4 97.0 84.8 68.6 59.8
E2E-MFD [72] NeurIPS’24 $\checkmark$ $\checkmark$ 31.3 37.1ms 76.1 91.9 61.1 91.8 91.3 97.2 84.9 69.5 60.5
E2E-OSDet - $\checkmark$ $\checkmark$ 27.5 20.9ms 77.7 90.7 64.3 91.8 92.1 97.8 85.7 70.3 61.4
Table 3: Ablation analysis of E2E-OSDet.
@@
(a) 83.4 66.4 58.0
(b) $\checkmark$ 84.5 67.1 58.9
(c) $\checkmark$ 84.9 68.6 59.6
(d) $\checkmark$ 84.6 68.9 60.0
(e) $\checkmark$ $\checkmark$ 85.5 69.9 61.0
(f) $\checkmark$ $\checkmark$ 84.9 69.5 60.7
(g) $\checkmark$ $\checkmark$ 85.2 70.0 60.9
(h) $\checkmark$ $\checkmark$ $\checkmark$ 85.7 70.3 61.4
Ablation Experiment. As shown in Table 3(a), we adopt CSSA [2] as our baseline model, which serves as the foundation for independently evaluating the contribution of each proposed module. The results (see Table 3 (a), (b), and (c)) indicate that all three modules lead to performance gains, among which the AFM module provides the most substantial improvement in detection accuracy. We further assess the interoperability of the modules. As shown in Table 3 (e), (f), and (g), the modules are compatible with one another, and their integration does not result in any performance degradation. Finally, we integrate all modules (see Table 3 (h)), resulting in additional gains in overall model performance.
@@
Figure 7: Visualization results of handcrafted features. SSIM represents structural similarity.
Impact of Filter Augment and Scanning Mechanisms. To assess the effectiveness of the filter-augmented inputs, we tested a range of conventional feature descriptors. As shown in Figure 6 (a) and Figure 7, the results indicate that incorporating handcrafted features significantly enhances detection performance. This finding suggests that mapping pixels into a handcrafted feature space effectively reduces the distribution gap between optical and SAR data. Notably, the Grad feature yields superior performance due to its ability to narrow domain differences and capture multi-scale information. We further evaluate the impact of the CMIM module under different scanning mechanisms [15]. As shown in Figure 6 (b), CMIM consistently enhances performance across all scanning strategies. In addition, we analyze the effects of the proportion of handcrafted feature input and the use of area attention. As shown in Figure 6(c) and (d), these hyperparameters exhibit marginal effects on performance and do not substantially influence the final results.
4.3 Qualitative Evaluation
We present visual comparisons of detection results from various multi-source methods. As shown in Figure 8, the proposed E2E-OSDet consistently outperforms other methods in complex scenes. In particular, in regions with limited optical information, such as cloud-covered, snow-covered, low-light, or low-resolution areas, the incorporation of SAR data significantly enhances target localization accuracy. These experimental results demonstrate the proposed dataset’s effectiveness in evaluating the robustness and adaptability of detection models across diverse scenarios. Additional qualitative results, including detection outputs and heatmaps, are provided in Appendix D.
5 Discussion
Summary. This work addresses the lack of publicly available, high-quality datasets for optical–SAR fusion object detection by introducing the M4-SAR dataset, the first large-scale dataset specifically tailored for this task. Furthermore, we propose E2E-OSDet, a novel detection framework that demonstrates strong robustness across diverse and challenging scenarios.
Limitations Analysis and Future Work. Despite the significant advantages of M4-SAR in terms of scale, diversity, and scene coverage, there remain several limitations that merit further exploration: First, a temporal gap exists between the acquisition of optical and SAR images, particularly affecting a minority of dynamic targets by introducing spatial inconsistencies. Future work could explore robust detection methods to handle temporal misalignment [68; 3; 10]. Second, although a semi-supervised annotation strategy is employed, SAR labels still rely on guidance from optical images, which can be unreliable under occlusion or in the presence of blurred boundaries. To address this, future research could introduce strategies for modeling missing modalities [57; 34] to enhance the model’s adaptability to weak or absent inputs. Third, the current dataset focuses on six representative static object categories and lacks more dynamic or fine-grained object categories. Expanding the object taxonomy and annotation granularity would benefit a wider range of remote sensing applications [58; 47]. Finally, although E2E-OSDet offers a strong performance baseline, its computational complexity hinders its deployment on resource-constrained edge devices. Future efforts could focus on developing lightweight multi-source detection frameworks (e.g., DecoupleNet [37], SAMNet [30]) to maintain high accuracy while improving inference efficiency.
Usefulness of M4-SAR. In benchmark evaluations, our proposed M4-SAR has demonstrated its effectiveness in leveraging the complementary characteristics of optical and SAR images to improve object detection accuracy. We believe that this dataset is not only suitable for multi-source fusion object detection but is also broadly applicable to cross-source reconstruction, multi-source alignment, multi-source image dehazing, and other related tasks.
@@ 5 Discussion
Figure 8: Qualitative comparison between the proposed E2E-OSDet and six fusion detection methods.
6 Conclusion

### v2

@@ IV-B3 Cross-modal Mamba Interaction Module
R-FCOS [43] $\checkmark$ 31.9M 17.6ms 32.8 79.8 29.7 67.0 78.9 54.8 57.2 26.4 30.2
R-ATSS [65] $\checkmark$ 36.0M 19.1ms 33.2 80.0 29.3 66.6 78.3 54.8 57.0 27.1 30.0
YOLOv5 [16] $\checkmark$ 8.10M 7.20ms 61.4 88.0 56.6 76.4 91.3 91.8 77.6 59.8 52.9
Roi-Trans [33] $\checkmark$ 273M 49.3ms 17.2 80.7 9.1 55.2 80.5 25.6 44.2 27.6 27.1
O-RepPoint [22] $\checkmark$ 36.6M 16.1ms 63.3 90.2 55.8 81.7 88.2 90.1 77.1 58.3 53.9
RTMDet [32] $\checkmark$ 8.86M 11.1ms 34.0 79.8 30.3 61.1 79.1 55.9 56.7 26.5 29.2
YOLOv8 [16] $\checkmark$ 10.0M 6.70ms 63.6 88.6 57.9 80.0 91.3 92.5 79.0 61.6 54.4
PSC [58] $\checkmark$ 31.9M 18.3ms 30.7 79.7 25.9 57.4 78.1 49.4 53.5 26.2 28.5
KFIoU [57] $\checkmark$ 36.2M 66.2ms 16.9 79.4 9.1 50.9 80.4 25.2 43.6 19.4 22.4
YOLOv6 [17] $\checkmark$ 16.0M 6.10ms 57.6 82.1 52.7 67.6 89.4 90.4 73.3 64.9 48.8
YOLOv9 [50] $\checkmark$ 6.40M 14.4ms 61.9 87.7 55.3 74.4 90.0 92.4 76.9 58.9 52.5
YOLOv10 [46] $\checkmark$ 7.50M 7.80ms 65.2 89.9 58.8 82.3 90.8 93.5 80.1 62.8 55.7
YOLOv11 [16] $\checkmark$ 9.74M 6.50ms 64.3 90.0 57.8 80.7 90.9 91.5 79.2 60.5 54.3
LSKNet [24] $\checkmark$ 21.8M 20.4ms 63.3 90.2 55.8 81.7 88.2 90.1 77.1 58.3 53.9
YOLOv12 [42] $\checkmark$ 9.36M 13.4ms 62.7 88.6 57.1 80.1 91.0 92.7 78.7 61.1 54.0
R-FCOS [43] $\checkmark$ 31.9M 17.6ms 23.9 80.1 27.9 60.1 75.8 43.5 51.9 26.4 28.3
R-ATSS [65] $\checkmark$ 36.0M 19.1ms 25.3 80.3 26.8 62.3 77.0 43.9 52.6 24.9 27.8
YOLOv5 [16] $\checkmark$ 8.10M 7.20ms 40.5 85.3 41.0 64.3 85.5 89.9 67.7 44.0 41.3
Roi-Trans [33] $\checkmark$ 273M 49.3ms 9.1 80.8 9.1 35.6 79.0 18.0 38.6 25.2 24.3
O-RepPoint [22] $\checkmark$ 36.6M 16.1ms 26.2 78.6 21.9 53.7 59.9 48.8 48.2 13.9 20.4
RTMDet [32] $\checkmark$ 8.86M 11.1ms 16.2 71.0 21.1 39.3 68.9 39.6 42.7 20.9 22.3
YOLOv8 [16] $\checkmark$ 10.0M 6.70ms 44.9 85.8 42.5 67.9 87.5 89.2 69.6 46.1 43.2
PSC [58] $\checkmark$ 31.9M 18.3ms 22.3 80.0 20.6 51.9 75.7 39.4 48.3 24.5 26.5
KFIoU [57] $\checkmark$ 36.2M 66.2ms 9.1 72.1 9.1 35.3 78.0 22.7 37.7 17.4 19.8
YOLOv6 [17] $\checkmark$ 16.0M 6.10ms 36.5 79.4 32.8 53.5 85.7 86.7 62.4 37.6 36.5
YOLOv9 [50] $\checkmark$ 6.40M 14.4ms 43.9 84.8 38.4 61.5 87.9 88.8 67.6 44.2 41.4
YOLOv10 [46] $\checkmark$ 7.50M 7.80ms 49.8 87.0 45.5 73.4 87.9 90.4 72.3 50.8 46.2
YOLOv11 [16] $\checkmark$ 9.74M 6.50ms 46.9 87.2 44.6 71.1 88.8 89.2 71.3 48.6 44.7
LSKNet [24] $\checkmark$ 21.8M 20.4ms 33.4 80.2 28.4 69.3 76.8 51.2 56.6 27.1 30.4
YOLOv12 [42] $\checkmark$ 9.36M 13.0ms 46.4 87.7 43.6 70.8 90.0 90.7 71.5 49.2 45.4
CFT [36] $\checkmark$ $\checkmark$ 53.8M 40.6ms 75.8 92.5 61.3 91.6 90.3 96.3 84.6 68.9 59.9
COMO [27] $\checkmark$ $\checkmark$ 22.6M 37.1ms 72.3 90.8 60.4 88.7 90.5 96.3 83.2 66.3 57.9
CLANet [14] $\checkmark$ $\checkmark$ 48.2M 29.1ms 74.8 92.2 60.7 91.3 91.6 97.2 84.6 68.5 59.6
CSSA [2] $\checkmark$ $\checkmark$ 13.5M 12.3ms 73.3 91.7 59.3 88.9 91.6 95.8 83.4 66.4 58.0
CMADet [39] $\checkmark$ $\checkmark$ 41.5M 46.7ms 70.9 90.7 52.0 86.4 91.7 97.1 81.5 63.5 55.7
ICAFusion [38] $\checkmark$ $\checkmark$ 29.0M 23.6ms 74.7 91.9 60.9 91.0 91.8 96.7 84.5 67.3 58.8
MMIDet [60] $\checkmark$ $\checkmark$ 53.8M 41.9ms 74.9 92.6 61.1 91.7 91.4 97.0 84.8 68.6 59.8
E2E-MFD [63] $\checkmark$ $\checkmark$ 31.3M 37.1ms 76.1 91.9 61.1 91.8 91.3 97.2 84.9 69.5 60.5
E2E-OSDet (Ours) $\checkmark$ $\checkmark$ 24.7M 20.9ms 77.7 90.7 64.3 91.8 92.1 97.8 85.7 70.3 61.4
IV-B4 Area-attention Fusion Module
To enhance the discriminative capacity of fused optical-SAR features across spatially diverse regions, we propose the Area-attention Fusion Module (AFM), a novel mechanism that emphasizes salient local structures and amplifies responses in critical target regions. AFM exploits the complementary characteristics of optical and SAR modalities, enabling robust and adaptive feature integration for multi-source object detection. Inspired by the efficient spatial modeling techniques in [42, 8], AFM partitions the input feature map of resolution $(H,W)$ into $k$ non-overlapping blocks of size $(H/k,W)$ or $(H,W/k)$, without requiring explicit window segmentation. This lightweight partitioning is implemented via a simple reshaping operation, preserving spatial continuity while substantially improving computational efficiency. The fused feature representation is computed as:
$F^{fus}_{j}=\mathbb{AFM}(F^{\prime}_{Oj},F^{\prime}_{Sj}),\quad j\in\{3,4,5\},$ (10)
where $F^{\prime}_{Oj}$ and $F^{\prime}_{Sj}$ denote the optical and SAR features. By adaptively weighting localized features, AFM improves the model’s capability to detect targets in complex remote sensing scenes characterized by varying object scales and cluttered backgrounds. As illustrated in Fig. 4 (c), AFM achieves efficient and effective cross-modal fusion, surpassing conventional spatial fusion strategies that often overlook regional saliency and incur higher computational overhead.
IV-C Loss Function
$\mathcal{L}_{total}=\mathcal{L}_{reg}+\mathcal{L}_{dfl}+\mathcal{L}_{cls},$ (11)
where $\mathcal{L}_{reg}$ denotes the bounding box regression loss, $\mathcal{L}_{dfl}$ represents the distribution focus loss, and $\mathcal{L}_{cls}$ corresponds to the classification loss. This formulation balances localization accuracy and classification confidence, ensuring end-to-end optimization of the optical-SAR fusion detection pipeline.
V M4-SAR Benchmark
V-A Setups and Implementation Details
@@
We also evaluate single-source baselines implemented via the MMRotate toolkit [35], including Rotated FCOS [43], Rotated ATSS [65], Oriented R-CNN [56], Oriented RepPoints [22], RTMDet [32], PSC [58], and LSKNet [24]. These methods use a ResNet-50 backbone [13] pre-trained on ImageNet, and are optimized using AdamW [30] (learning rate: 2e-4, weight decay: 0.05) for 72 epochs with a batch size of 32 and the same $\times512$ input resolution. During evaluation, we adopted the standard COCO evaluation protocol to measure Average Precision (AP), including $AP_{50}$, $AP_{75}$, and $mAP$.
@@
TABLE III: Ablation analysis of E2E-OSDet on M4-SAR dataset.
[…]
(a) 14.17M baseline 83.4 66.4 58.0
(b) $\checkmark$ 14.17M +0M 84.5 67.1 58.9
(c) $\checkmark$ 19.01M +5.68M 84.9 68.6 59.6
(d) $\checkmark$ 19.84M +4.84M 84.6 68.9 60.0
(e) $\checkmark$ $\checkmark$ 19.84M +4.84M 85.5 69.9 61.0
(f) $\checkmark$ $\checkmark$ 19.01M +5.68M 84.9 69.5 60.7
(g) $\checkmark$ $\checkmark$ 24.69M +10.52M 85.2 70.0 60.9
(h) $\checkmark$ $\checkmark$ $\checkmark$ 24.69M +10.52M 85.7 70.3 61.4
V-B Main Results
We first compared multi-source fusion methods (such as CFT, CLANet, CSSA, CMADet, ICAFusion, MMIDet, and E2E-MFD) with single-source detection methods (such as R-FCOS, R-ATSS, Roi-Trans, O-RepPoint, RTMDet, and the YOLO series). Experimental results (see Table II) indicate that detection accuracy using SAR images alone is significantly lower than when using optical images alone. This gap primarily stems from the unique imaging mechanism of SAR: even when targets are visible in the image, SAR struggles to accurately estimate the shape of target bounding boxes.
In contrast, all multi-source fusion methods outperform their corresponding single-modal baselines, clearly demonstrating that cross-modal fusion effectively leverages the complementarity between SAR and optical data to significantly reduce false positives and enhance detection accuracy. Among these, our proposed E2E-OSDet achieves the highest performance across all fusion methods, with a $mAP$ of 61.4%, fully validating the importance of mitigating domain differences in cross-modal detection. Furthermore, E2E-OSDet exhibits high efficiency with only 24.7M parameters and achieves an inference speed of 20.9ms. We also evaluated its generalization capability across different detection frameworks (see Fig. 6 (e)), demonstrating strong transferability and robustness.
Figure 7: Visualization results of handcrafted features.
@@ V-B Main Results
Figure 8: Qualitative comparison between the proposed E2E-OSDet and six fusion detection methods on the proposed M4-SAR dataset.
V-C Ablation Experiment
As shown in Table III (a), we adopt CSSA [2] as our baseline model, which serves as the foundation for independently evaluating the contribution of each proposed module. The results (see Table III (a), (b), and (c)) indicate that all three modules lead to performance gains, among which the AFM module provides the most substantial improvement in detection accuracy. We further assess the interoperability of the modules. As shown in Table III (e), (f), and (g), the modules are compatible with one another, and their integration does not result in any performance degradation. Finally, we integrate all modules (see Table III (h)), resulting in additional gains in overall model performance.
V-C1 Impact of Filter Augment
To assess the effectiveness of the FAM, we tested a range of conventional feature descriptors. As shown in Fig. 6 (a) and Fig. 7, the results indicate that incorporating handcrafted features significantly enhances detection performance. These results mapping pixels into a handcrafted feature space effectively reduces the distribution gap between optical and SAR data. Notably, the Grad feature yields superior performance due to its ability to narrow domain differences and capture multi-scale information.
V-C2 Impact of Scanning Mechanisms
We further investigate the effectiveness of the CMIM module under different scanning mechanisms [11]. As illustrated in Fig. 6 (b), CMIM consistently enhances detection accuracy regardless of the underlying scanning strategy. This observation highlights the robustness and adaptability of CMIM, suggesting that the proposed module can serve as a versatile component that generalizes well across diverse architectural settings rather than being restricted to a particular scanning design.
V-C3 Impact of Feature Proportion and Area Size
In addition, we investigate the influence of two key design choices: the proportion of handcrafted feature input and the adoption of area attention. As shown in Fig. 6 (c) and (d), varying these hyperparameters leads to only marginal performance differences, indicating that the overall framework is relatively insensitive to such configurations. This stability suggests that our method achieves consistent performance without requiring delicate hyperparameter tuning, further underscoring its practicality in real-world deployment.
V-D Qualitative Evaluation
We present visual comparisons of detection results from various multi-source methods. As shown in Fig. 8, the proposed E2E-OSDet consistently outperforms other methods in complex scenes. In particular, in regions with limited optical information, such as cloud-covered, snow-covered, low-light, or low-resolution areas, the incorporation of SAR data significantly enhances target localization accuracy. As illustrated in Fig. 9, our proposed E2E-OSDet produces highly concentrated and spatially aligned attention maps, accurately focusing on target regions while suppressing irrelevant background activations. These experimental results demonstrate the proposed dataset’s effectiveness in evaluating the robustness and adaptability of detection models across diverse scenarios.
Figure 9: Grad-CAM [37] heatmaps of the proposed E2E-OSDet and six fusion detection methods on the proposed M4-SAR dataset.
VI Discussions
VI-A Expanded Application of E2E-OSDet
To further assess the generalization capability of the E2E-OSDet framework, we conducted additional experiments on the OGSOD-1.0 [49] and OGSOD-2.0 [48] datasets (300 epochs, input resolution of $\times256$). As shown in Table IV, E2E-OSDet attains the highest performance on OGSOD-1.0, achieving an $mAP$ of 67.3% and surpassing representative multimodal methods, including CLANet, ICAFusion, and MMIDet. These results demonstrate the robustness and generalizability of E2E-OSDet across various benchmarks.
VI-B Limitations Analysis of M4-SAR
Despite the substantial advantages of M4-SAR in terms of scale, diversity, and scene coverage, certain limitations persist that require further exploration. First, a temporal misalignment often occurs between the acquisition of optical and SAR images, which may cause spatial inconsistencies for certain dynamic targets. Second, the dataset is currently limited to six representative static categories and does not yet include more dynamic or fine-grained object classes. Third, although E2E-OSDet provides a robust performance baseline, its computational demands still present challenges for deployment on resource-constrained edge devices.
VI-C Future Work
M4-SAR offers a high-quality benchmark for optical-SAR fusion detection, with E2E-OSDet delivering superior performance both within this dataset and in cross-dataset evaluations. Future research could investigate the following directions: 1. Broadening the dataset categories to include dynamic and fine-grained targets. 2. Developing lightweight network architectures for edge deployment compatibility. 3. Implementing temporal alignment and multimodal temporal modeling to address cross-modal acquisition time disparities.
TABLE IV: Quantitative results on the OGSOD-1.0 and OGSOD-2.0.
  Method O S OGSOD-1.0[49] OGSOD-2.0[48]
$\textbf{AP}_{50}$ $\textbf{AP}_{75}$ mAP $\textbf{AP}_{50}$ $\textbf{AP}_{75}$ mAP
YOLOv11 [16] $\checkmark$ 90.8 66.0 61.5 88.2 62.6 58.5
YOLOv11 [16] $\checkmark$ 76.0 48.0 47.8 73.8 46.2 45.9
CFT [36] $\checkmark$ $\checkmark$ 92.9 72.1 65.3 90.8 67.1 61.4
CLANet [14] $\checkmark$ $\checkmark$ 93.5 72.9 66.3 91.6 68.7 63.0
CSSA [2] $\checkmark$ $\checkmark$ 92.1 69.3 63.3 89.7 64.2 59.8
CMADet [39] $\checkmark$ $\checkmark$ 91.3 67.6 62.9 88.1 60.9 57.7
ICAFusion [38] $\checkmark$ $\checkmark$ 93.5 72.7 66.1 90.7 68.2 62.2
MMIDet [60] $\checkmark$ $\checkmark$ 93.5 73.5 66.7 91.6 68.6 63.0
E2E-OSDet $\checkmark$ $\checkmark$ 93.6 74.3 67.3 92.3 71.8 65.1
VII Conclusion